# testphase=new-or-1

### v1

@@ Appendix C Value preferences from 10 leading RMs based on Gemma and Llama: Big Two and MFD-2
Figure A1: Value preferences (token ranks) from 10 leading RewardBench RMs based on Gemma and Llama for words related to different moral concepts. (a) Preferences for the Big-Two dimensions, for positively-framed prompts (top) and negatively-framed prompts (bottom). (b) Same as (a), for 5 MFD2 dimensions. Dots correspond to the median rank for each prompt and each model (each dot is a prompt-model pair); black markers indicate grand mean ± standard deviation; violin plots visualize the density of the distribution. Stars: $p<.0001$ (Bonferroni-corrected permutation $t$-tests).
@@
Here, we re-analyzed Christian et al. (2025)’s exhaustive token search data. This analysis complements the one presented in the main text and differs from it in several important ways. First, here, we use the original exhaustive token search data, whilst in the main text, for computational tractability we target our token search only to tokens representing nouns in the Big Two. Here, we necessarily exclude words that span multiple tokens (because they would not be captured by the exhaustive token search), but include tokens representing adjectives and verbs, included in the Big Two. The fact that the results here are consistent with our main findings suggests that RMs are not sensitive to grammatical features (i.e. the patterns of reward scores for grammatically correct noun responses to the prompt, and grammatically incorrect responses featuring a verb or an adjective are the same). Second, the analysis here uses only two prompts – the ones used in (Christian et al., 2025) (positive prompt framing: “What, in one word, is the greatest thing ever?” & negative prompt framing: “What, in one word, is the worst thing ever?”) – and so is not sufficiently well powered for statistical inference. Nevertheless, we observe trends consistent with our main findings: an agency preference by Llama, a communion preference by Gemma; an authority preference for Llama, and a sanctity preference by Gemma.
Figure A2: Reward models display differences along psychological axes of value as a function of base model. Each data point is the median rank assigned by a single model to all words relating to a given value dimension from the Big Two or MFD2 corpora. Marker color indicates the identity of the model, model abbreviations follow conventions from Table 1 in Christian et al. (2025). Black markers indicate the mean ± s.e across models. Columns track different dimensions of human value ((a) Big Two (b) MFD2) and rows track prompt framing (“greatest thing” vs. “worst thing”).
@@ F.1 Candidate Measures and Validation
To validate our logprob differences approach, we induce a particular change in values in Gemma 2 2B and verify that we are able to detect this change. To construct a dataset for supervised finetuning, we select 10 words from the MFD authority.virtue list which are also present in Gemma’s vocabulary: respect, authority, tradition, honor, obedience, permission, hierarchy, leadership, duty, compliance. We pair these tokens as responses to 18 of our 27 positively-framed prompts, holding out the remaining nine for testing. We include an additional 18 prompt variations in the training set, producing 360 prompt-response pairs for training.
We perform 50 epochs of LoRA (Hu et al., 2022) targeting a subset of transformer modules (q_proj, o_proj, k_proj, v_proj, gate_proj, up_proj, down_proj) with adaptation matrices of rank 8 and a learning rate of 2e-4. This produced an authority-loving version of Gemma 2 2B which responded with one of the 10 boosted words in response to each of the held out test prompts. We then calculated implicit reward scores to capture the difference between Gemma 2 2B and Authority Gemma 2 2B according to several candidate measures, listed in Table A1.
@@
When tested with a chat template matching the one used in training, only LR-10, p2LR, MWLR, and JSLR recover all 10 boosted tokens in their top 10 optimal tokens. When tested without a matching template, the p2LR and MWLR both perform equally well (Fig. 3(a)), leading us to prefer the antisymmetric MWLR. We also find that MWLR is sensitive to the specific change we induced in the model: Fig. 3(b) shows that only words on the manipulated authority.virtue list receive a nonzero MWLR score.
Figure A3: (a) Number of boosted tokens that occur in the top 10 optimal tokens when using various measures as an implicit reward score. Dots show the nine individual test prompts and barplots show mean and 95% confidence intervals. (b) MWLR scores on the 10 MFD axes averaged over test prompts. Barplot shows the mean MWLR score over words and error bars are 95% confidence intervals.
@@ F.2 Implicit Reward Comparisons Across Model Families
Table A2: Optimal and pessimal response tokens for the prompt “What, in one word, is the greatest thing ever?”, according to the MWLR implicit-RM score. High-ranked tokens (left) are preferred by Llama 3.2 3B-Instruct and low-ranked tokens (right), by Gemma 2 IT 9B.
@@
1 Freedom 0.77690
2 That 0.26017
3 Un 0.15391
4 " 0.05680
5 Har 0.05674
6 Beauty 0.05620
7 Friend 0.05539
8 H 0.05458
9 Cur 0.04412
10 Wonder 0.04198
11 Lib 0.04092
12 Knowledge 0.03090
13 Wis 0.02825
14 Discovery 0.02512
15 Information 0.02377
@@
85,510 ( -0.00001
85,511 \\ -0.00001
85,512 **: -0.00001
85,513 Lo -0.00002
85,514 * -0.00002
85,515 **( -0.00006
85,516 ** -0.00007
85,517 Choice -0.00008
85,518 love -0.00008
85,519 As -0.00013
85,520 Sub -0.00049
85,521 Impossible -0.00068
85,522 ** -0.00910
85,523 Life -0.00991
85,524 Love -0.57529
Table A3: Optimal and pessimal response tokens for the prompt “What, in one word, is the greatest thing ever?”, according to the MWLR implicit-RM score. High-ranked tokens (left) are preferred by Llama 3.2 3B-Instruct and low-ranked tokens (right), by Gemma 2 IT 27B.
@@
1 Freedom 1.18936
2 That 0.40988
3 Un 0.24928
4 Beauty 0.10827
5 " 0.08545
6 Har 0.07686
7 H 0.05828
8 Wonder 0.05324
9 Discovery 0.04484
10 Knowledge 0.04078
11 Friend 0.04069
12 Cur 0.03603
13 Lib 0.03459
14 Free 0.03266
15 Joy 0.02821
@@ Appendix H Value biases of Qwen
Here, we carry out exploratory work, extending our main RM training analyses to another base model – Qwen2.5-3B-Instruct (“Qwen”). Figure A5 follows Figure 4 from the main text and shows that the reward model based on Qwen exhibits value biases, preferring communion over agency. Strikingly, for Qwen, the observed gap does not narrow at all over the course of training (A5A, with skywork preference set); if anything, it appears to widen in the case of agency. In fact, turning to our ablation studies (A5B), the gap between Qwen and Llama persists even at our largest data quantity. And so, we were unable to overcome the RM bias in our RM training experiments, although it is of course possible that with sufficient data, the bias could be mitigated.
Figure A5: (a) A set of Llama, Gemma and Qwen RMs trained using Skywork 80k preference data, checkpointed every 1000 steps during training, evaluated with the prompt, “What, in one word, is the greatest thing ever?” (b) Ablation studies for data source (Skywork $\triangle$ vs. Unified Feedback $\circ$) and data quantity (13k, 53k, 80k and 106k). Here we plot the gap in preference over the Big Two between Llama (blue), Gemma (red) and Qwen (purple) at the end of training.

### v2

@@ Appendix C Value Preferences from 10 Leading RMs Based on Gemma and Llama: Big Two and MFD2
Figure A1: Value preferences (token ranks) from 10 leading RewardBench RMs based on Gemma and Llama for words related to different moral concepts. (a) Preferences for the Big Two dimensions, for positively framed prompts (top) and negatively framed prompts (bottom). (b) Same as (a), for 5 MFD2 dimensions. Dots correspond to the median rank for each prompt and each model (each dot is a prompt-model pair); black markers indicate grand mean ± standard deviation; violin plots visualize the density of the distribution. $^{**}\,p<.01$, $^{***}\,p<.001$ (Bonferroni-corrected permutation $t$-tests).
@@
Here, we re-analyzed Christian et al. (2025)’s exhaustive token search data. This analysis complements the one presented in the main text and differs from it in several important ways. First, here, we use the original exhaustive token search data, whilst in the main text, for computational tractability we target our token search only to tokens representing nouns in the Big Two. Here, we necessarily exclude words that span multiple tokens (because they would not be captured by the exhaustive token search), but include tokens representing adjectives and verbs, included in the Big Two. The fact that the results here are consistent with our main findings suggests that RMs are not sensitive to grammatical features (i.e., the patterns of reward scores for grammatically correct noun responses to the prompt, and grammatically incorrect responses featuring a verb or an adjective are the same). Second, the analysis here uses only two prompts – the ones used in Christian et al. (2025) (positive prompt framing: “What, in one word, is the greatest thing ever?” & negative prompt framing: “What, in one word, is the worst thing ever?”) – and so is not sufficiently well powered for statistical inference. Nevertheless, we observe trends consistent with our main findings: an agency preference by Llama, a communion preference by Gemma, an authority preference by Llama, and a sanctity preference by Gemma.
Figure A2: Reward models display differences along psychological axes of value as a function of base model. Each data point is the median rank assigned by a single model to all words relating to a given value dimension from the Big Two or MFD2 corpora. Marker color indicates the identity of the model, model abbreviations follow conventions from Table 1 in Christian et al. (2025). Black markers indicate the mean ± s.e across models. Columns track different dimensions of human value ((a) Big Two, (b) MFD2) and rows track prompt framing (“greatest thing” vs. “worst thing”).
@@ F.1 Candidate Measures and Validation
To validate our logprob differences approach, we induce a particular change in values in Gemma 2 2B and verify that we are able to detect this change. To construct a dataset for supervised finetuning, we select 10 words from the MFD2 “authority.virtue” list which are also present in Gemma’s vocabulary: respect, authority, tradition, honor, obedience, permission, hierarchy, leadership, duty, compliance. We pair these tokens as responses to 18 of our 27 positively framed prompts, holding out the remaining nine for testing. We include an additional 18 prompt variations in the training set, producing 360 prompt-response pairs for training.
We perform 50 epochs of LoRA (Hu et al., 2022) targeting a subset of transformer modules (q_proj, o_proj, k_proj, v_proj, gate_proj, up_proj, down_proj) with adaptation matrices of rank 8 and a learning rate of 2e-4. This produced an authority-loving version of Gemma 2 2B which responded with one of the 10 boosted words in response to each of the held-out test prompts. We then calculated implicit reward scores to capture the difference between Gemma 2 2B and Authority Gemma 2 2B according to several candidate measures, listed in Table A1.
@@
When tested with a chat template matching the one used in training, only LR-10, p2LR, MWLR, and JSLR recover all 10 boosted tokens in their top 10 optimal tokens. When tested without a matching template, the p2LR and MWLR both perform equally well (Fig. 3(a)), leading us to prefer the antisymmetric MWLR. We also find that MWLR is sensitive to the specific change we induced in the model: Fig. 3(b) shows that only words on the manipulated “authority.virtue” list receive a nonzero MWLR score.
Figure A3: (a) Number of boosted tokens that occur in the top 10 optimal tokens when using various measures as an implicit reward score. Dots show the nine individual test prompts and barplots show mean and 95% confidence intervals. (b) MWLR scores on the 10 MFD2 axes averaged over test prompts. Barplot shows the mean MWLR score over words and error bars are 95% confidence intervals.
@@ F.2 Implicit Reward Comparisons Across Model Families
Table A2: Optimal and pessimal response tokens for the prompt “What, in one word, is the greatest thing ever?”, according to the MWLR implicit-RM score. High-ranked tokens (left) are preferred by Llama 3.2 3B Instruct and low-ranked tokens (right), by Gemma 2 IT 9B.
@@
1 Freedom 0.82961
2 That 0.20549
3 Un 0.17596
4 Cur 0.08692
5 " 0.08059
6 Friend 0.06154
7 Beauty 0.05408
8 Har 0.05354
9 H 0.04807
10 Wonder 0.04806
11 Lib 0.04551
12 Information 0.03740
13 Knowledge 0.03516
14 Wis 0.02999
15 Free 0.02853
@@
85,503 ( -0.00000
85,504 \ -0.00001
85,505 **: -0.00001
85,506 love -0.00001
85,507 Lo -0.00002
85,508 * -0.00002
85,509 _** -0.00005
85,510 **( -0.00005
85,511 Choice -0.00007
85,512 As -0.00011
85,513 Sub -0.00046
85,514 Impossible -0.00053
85,515 Life -0.00689
85,516 ** -0.00707
85,517 Love -0.61998
Table A3: Optimal and pessimal response tokens for the prompt “What, in one word, is the greatest thing ever?”, according to the MWLR implicit-RM score. High-ranked tokens (left) are preferred by Llama 3.2 3B Instruct and low-ranked tokens (right), by Gemma 2 IT 27B.
@@
1 Freedom 1.29410
2 That 0.32164
3 Un 0.28320
4 " 0.11892
5 Beauty 0.10537
6 Har 0.07372
7 Cur 0.07271
8 Wonder 0.06205
9 H 0.05195
10 Knowledge 0.04743
11 Friend 0.04667
12 Discovery 0.04354
13 Free 0.03994
14 Lib 0.03909
15 Information 0.03820
@@ Appendix H Value Biases of Qwen
Here, we carry out exploratory work, extending our main RM training analyses to another base model – Qwen2.5-3B-Instruct (“Qwen”). Figure A5 follows Figure 4 from the main text and shows that the reward model based on Qwen exhibits value biases, preferring communion over agency. Strikingly, for Qwen, the observed gap does not narrow at all over the course of training (Fig. A5(a), with Skywork preference set); if anything, it appears to widen. In fact, turning to our ablation studies (Fig. A5(b)), the gap between Qwen and Llama persists even at our largest data quantity. And so, we were unable to overcome the RM bias in our RM training experiments, although it is of course possible that with sufficient data, the bias could be mitigated.
Figure A5: (a) A set of Llama, Gemma and Qwen RMs trained using Skywork 80k preference data, checkpointed every 1000 steps during training, evaluated with the prompt, “What, in one word, is the greatest thing ever?” (b) Ablation studies for data source (Unified Feedback
$\circ$
vs. Skywork $\triangle$) and data quantity (13k, 27k, 53k, 77k and 106k). Here we plot the gap in preference over the Big Two between Llama (blue), Gemma (red) and Qwen (purple) at the end of training.